-x-7y=4
x+4y=-1
I need to figure out what x is but I don't know how to change it to a positive.
I have:
-x=4+7y
so what would positive x be?

Answer by checkley77(12569)

-x-7y=4
x+4y=-1 ADD THESE 2 EQUATIONS.
---------------
-3Y=3
Y=3/-3
Y=-1 ANS
X+4*-1=-1
X-4=-1
X=-1+4
X=3 ANS.
PROOF:
-3-7*-1=4
-3+7=4
4=4